Google:

Part A: If (7^2)^x = 1, what is the value of x? Explain your answer. (5 points) Part B: If (7^0)^x = 1, what are the possible values of x? Explain your answer. (5 points)

iGreen:

Anything to the power of 0 is 1.

iGreen:

1 to the power of anything is 1

MARC:

Part A: \((7^2)^x=1\) \(49^x=1\) \(49^x=49^0\) \(x=?\)

MARC:

If not mistaken, Part B: \((7^0)^x=1\) \(1^x=1\) \(1^x=1^0\) or \(1^x=1^1\) or \(1^x=1^{1000}\)... which means that the value of x is infinity.